Assume Anderson’s General Store bought, on credit, a truckload of merchandise from American Wholesaling costing $23,900. If Anderson’s paid National Trucking $740 cash for transportation, immediately returned goods to American Wholesaling costing $1,200, and then paid American Wholesaling within the 2/30, n/60 purchase discount period.

How much did this inventory cost Anderson’s? Assume Anderson’s uses a perpetual inventory system.

Inventory Cost $
Answer #1

Answer) $22986

Explanations:

Cost of Inventory = Cost of goods purchased + Transportation cost - Cost of goods returned - Discount received for prompt payment

= 23900 + 740 - 1200 - 454

= 22986

*Discount received = (23900 - 1200) x 2% = 454

*Since the company uses perpetual inventory system, transportation cost is added to inventory and discount received is deducted from inventory.